# Integrative Learning of Intensity Fluctuations of Quantum Dots under Excitation via a Tailored Mixture Hidden Markov Model

### v1

@@ 5.2 Results from 3-Cluster Model
for $h=1,\ldots,3$. As seen from both Figure 5 and the estimated parameters in Table 4, the three states clearly correspond to relatively low, medium, and high intensity levels. The low-intensity state (State 1) accommodates intensities in the range of about 0.00 to 0.50, with mean 0.290 and relatively the highest variance. The medium-intensity state (State 2) mostly concentrates within the 0.50 to 0.80 spectrum, with mean 0.597 and a relatively small variance. The high-intensity state (State 3) covers the range of about 0.75 to 1.00, with mean 0.787 and relatively the smallest variance. The zero and one inflation rates are also reflected in Figure 5. The low-intensity state accommodates the zero intensities, while the high-intensity state encompasses all the intensities elevated to one.
@@
State 1 State 2 State 3
$\widehat{\epsilon}^{(0)}$ 0.025 $(1.163\times 10^{-3})$ 0.000 $(5.742\times 10^{-6})$ 0.000 $(1.145\times 10^{-13})$
$\widehat{\epsilon}^{(1)}$ 0.000 $(1.521\times 10^{-6})$ 0.001 $(4.091\times 10^{-3})$ 0.017 $(4.803\times 10^{-6})$
$\widehat{a}$ 2.195 $(1.900\times 10^{-2})$ 10.077 $(2.437\times 10^{-4})$ 11.658 $(2.656\times 10^{-3})$
$\widehat{b}$ 5.183 $(4.06\times 10^{-2})$ 6.805 $(6.239\times 10^{-5})$ 3.227 $(1.340\times 10^{-3})$
Mean 0.290 0.597 0.787
Variance 0.0266 0.0137 0.0113
@@ 5.3 State Transition Patterns and Clusters of Quantum Dots
$\widehat{\boldsymbol{\pi}}_{1}=\begin{bmatrix}0.213&0.443&0.344\end{bmatrix},\widehat{\boldsymbol{\pi}}_{2}=\begin{bmatrix}0.000&0.089&0.911\end{bmatrix},\widehat{\boldsymbol{\pi}}_{3}=\begin{bmatrix}0.106&0.459&0.435\end{bmatrix}.$
Based on the estimated MHMM-$\beta$ model and according to the Bayes rule, there are 28, 54, and 46 QDs in Clusters 1, 2, and 3, respectively. To visualize, we randomly pick 3 dots from each cluster and presented their intensity series along with their estimated state serires in Figures 6, 7, and 8, respectively.
@@
In Cluster 2, the QDs QDs exhibit extremely high probabilities of staying in the medium or high-intensity states, and more importantly, the probabilities of transiting into the low-intensity state are essentially zero. This results in a stationary distribution showing that these QDs spend most of their time in the high-intensity state with probability 91.1%, a much smaller amount of time in the medium-intensity state with probability 8.9%, and ultimately no time at all in the low-intensity state. In Figure 7, the three randomly selected QDs from this cluster all stayed in the high-intensity state throughout the continuous excitation.
@@ 6 Conclusion & Discussion
This study introduces a novel statistical approach to analyzing the intensity fluctuation patterns of colloidal quantum dots (QDs), a critical problem in the fields of chemistry and materials science. The main challenge lies in the inherent complexity of the data: the intensity measurements exhibit intricate patterns influenced by unobserved states and are subject to substantial variation across individual QDs. The proposed mixture hidden Markov model with an inflated Beta distribution (MHMM-$\beta$) addresses these challenges by simultaneously modeling the intensity fluctuations and clustering the QDs based on their transition dynamics. This integrative approach captures the shared structure among QDs while allowing for individual differences, providing a more comprehensive understanding of their individual and collective behaviors under continuous excitation.
The findings from our analysis highlight the method’s effectiveness and its ability to generate interpretable results that align with principles of chemistry and physics. By identifying three distinct clusters of QDs, each characterized by unique transition dynamics and stationary distributions, our approach provides chemists with valuable insights into the interplay between QD intensity states and their fluctuation styles. These results not only validate the utility of the proposed method but also demonstrate its potential to guide future research and applications of QDs.
Looking ahead, there are several promising directions for extending this work. One critical avenue is the joint analysis of photon lifetime and intensity. Recent chemical studies have emphasized the importance of understanding the relationship between photon lifetime, which reflects aspects of the QD microenvironment, and intensity fluctuations, which capture macroenvironmental behaviors. Developing methods that incorporate photon lifetime into the integrative analysis would enable us to reveal dependencies and connections between these two dimensions, offering a more holistic view of QD properties.
Further methodological advancements could focus on developing more efficient algorithms for computation and implementing more accurate methods for model selection, particularly for determining the optimal number of clusters. Relaxing the assumptions of shared state distributions across clusters could allow for greater flexibility in modeling. Additionally, exploring Bayesian frameworks for parameter estimation and uncertainty quantification may enhance the robustness and interpretability of the results. From a broader perspective, the proposed approach could be adapted to address other scientific problems involving complex temporal data with state-dependent dynamics. For example, similar methods could be applied in neuroimaging to study brain activity patterns or in ecology to analyze animal movement behaviors.
@@ Acknowledgment
Appendix A Computation Details
@@ A.1 Derivation of the likelihood functions
The forward probabilities are defined as
$\alpha_{i,t,k,h}=P[(x_{i,1},\ldots,x_{i,t})^{\rm T},s_{i,t,h}=1\mid z_{i,k}=1,\boldsymbol{\Theta}^{(j)}].$
Then $\alpha_{i,T,k,h}$, where $T$ denotes the last time point in a sequence, stands for the likelihood of hidden state sequence ending in state $h$. That is, all the hidden sequences of dot $i$ ending with state $h$ are considered. So, we can sum up $\alpha_{i,T,k,h}$ w.r.t. $h=\{1,\ldots,M\}$ to get the summation over all possible sequences in (3).
@@
$\begin{split}l(\boldsymbol{\Theta};\boldsymbol{X},\boldsymbol{S},\boldsymbol{Z})=&\sum_{i=1}^{N}\log\{\prod_{k=1}^{K}[\delta_{k}\boldsymbol{\pi}_{k,\boldsymbol{s}_{i,1}}f(x_{i,1}\mid\boldsymbol{s}_{i,1})\prod_{t=2}^{T}(\boldsymbol{\Pi}_{k})_{\boldsymbol{s}_{i,t-1},\boldsymbol{s}_{i,t}}f(x_{i,t}\mid\boldsymbol{s}_{i,t})]^{I(z_{i,k}=1)}\}\\ =&\sum_{i=1}^{N}\sum_{k=1}^{K}I(z_{i,k}=1)\log(\delta_{k})+\sum_{i=1}^{N}\sum_{k=1}^{K}I(z_{i,k}=1)\sum_{p=1}^{M}I(s_{i,1,p}=1)\log(\boldsymbol{\pi}_{k,p})\\ &+\sum_{i=1}^{N}\sum_{k=1}^{K}I(z_{i,k}=1)\log(\prod_{t=2}^{T}(\boldsymbol{\Pi}_{k})_{\boldsymbol{s}_{i,t-1},\boldsymbol{s}_{i,t}})+\sum_{i=1}^{N}\sum_{k=1}^{K}I(z_{i,k}=1)\log(\prod_{t=1}^{T}f(x_{i,t}\mid\boldsymbol{s}_{i,t}))\\ =&\sum_{i=1}^{N}\sum_{k=1}^{K}I(z_{i,k}=1)\log(\delta_{k})+\sum_{i=1}^{N}\sum_{k=1}^{K}I(z_{i,k}=1)\sum_{p=1}^{M}I(s_{i,1,p}=1)\log(\boldsymbol{\pi}_{k,p})\\ &+\sum_{i=1}^{N}\sum_{k=1}^{K}I(z_{i,k}=1)\{\sum_{t=2}^{T}\sum_{p=1}^{M}\sum_{q=1}^{M}I(s_{i,t-1,p}=1,s_{i,t,q}=1)log((\boldsymbol{\Pi}_{k})_{p,q}\}\\ &+\sum_{i=1}^{N}\sum_{k=1}^{K}I(z_{i,k}=1)\{\sum_{t=1}^{T}\sum_{p=1}^{M}I(s_{i,t,p}=1)\log(f(x_{i,t}\mid s_{i,t,p}=1))\}.\end{split}$ (7)
@@ A.2 E-Step
Let $\boldsymbol{\Theta}^{(j)}$ denote the set of parameter estimates from $j^{th}$ iteration. Let
@@
In (8), we rewrite the integration w.r.t. the joint distribution $(\boldsymbol{S},\boldsymbol{Z})$ into a double-expectation structure, which could simplify the computation. For each dot $i$, given the observed intensities and parameters $\boldsymbol{\Theta}^{(j)}$, we will present the formula for all unknown values in (5).
@@
$E[I(z_{i,k}=1)\mid\boldsymbol{x}_{i},\boldsymbol{\Theta}^{(j)}]$ $=$ $Pr(z_{i,k}=1\mid\boldsymbol{x}_{i},\boldsymbol{\Theta}^{(j)})$ (9)
$=$ $\frac{Pr(z_{i,k}=1,\boldsymbol{x}_{i}\mid\boldsymbol{\Theta}^{(j)})}{\sum_{k^{\prime}=1,\dots,K}Pr(z_{i,k^{\prime}}=1,\boldsymbol{x}_{i}\mid\boldsymbol{\Theta}^{(j)})}$
Within formula (9), the key part is $Pr(z_{i,k}=1,\boldsymbol{x}_{i}\mid\boldsymbol{\Theta}^{(j)})$. To calculate this probability, we adopt the idea of forward-backward algorithm in dynamic programming (Rabiner, 1989). That is
$\alpha_{i,t,k,h}=Pr((x_{i,1},\dots,x_{i,t})^{\rm T},s_{i,t,h}=1\mid z_{i,k}=1,\boldsymbol{\Theta}^{(j)})$ (10)
The value of $\boldsymbol{\alpha}_{i,t,k}$ can be determined through iterative computation (11).
@@
$\boldsymbol{\alpha_{i,t,k}}$ $=$ $(\boldsymbol{\Pi}_{k}^{\rm T}\cdot\boldsymbol{\alpha}_{i,t-1,k})\circ(f(x_{i,t})\mid s_{i,t,h}=1,\boldsymbol{\Theta}^{(j)})_{h=1,\dots,M}^{\rm T},\textnormal{when }t>1,$ (11)
@@
Using the forward probabilities at the last time point $T$, we have $Pr(z_{i,k}=1,\boldsymbol{x}_{i}\mid\boldsymbol{\Theta}^{(j)})=\lVert\boldsymbol{\alpha}_{i,T,k}\rVert_{1}=\sum_{h=1}^{M}\alpha_{i,T,k,h}$ and update (9) to (12):
$E[I(z_{i,k}=1)\mid\boldsymbol{x}_{i},\boldsymbol{\Theta}^{(j)}]$ $=$ $\frac{\lVert\boldsymbol{\alpha}_{i,T,k}\rVert_{1}}{\sum_{k^{\prime}=1,\dots,K}\lVert\boldsymbol{\alpha}_{i,T,k^{\prime}}\rVert_{1}}$ (12)
@@
Follow the notation from Rabiner (1989), we denote $E[I(z_{i,k}=1)\mid\boldsymbol{x}_{i},\boldsymbol{\Theta}^{(j)}]$ as $\tau_{i,k}(\boldsymbol{\Theta}^{(j)})$.
(ii) Compute $I(s_{i,t,p}=1)$ and $I(s_{i,1,p}=1)$ :
Only $E[I(s_{i,t,p}=1)]$ is solved and $E[I(s_{i,1,p}=1)]$ is treated as a special case of the previous formula. Continue with (5), after solving the inner expectation, the outer expectation is:
$E[I(s_{i,t,h}=1)\mid\boldsymbol{z}_{i},\boldsymbol{x}_{i},\boldsymbol{\Theta}^{(j)}]$ $=$ $Pr(s_{i,t,h}=1\mid\boldsymbol{z}_{i},\boldsymbol{x}_{i},\boldsymbol{\Theta}^{(j)})$ (13)
$=$ $\frac{Pr(s_{i,t,h}=1,\boldsymbol{x}_{i}\mid\boldsymbol{z}_{i},\boldsymbol{\Theta}^{(j)})}{\sum_{h^{\prime}=1,\dots,M}Pr(s_{i,t,h^{\prime}}=1,\boldsymbol{x}_{i}\mid\boldsymbol{z}_{i},\boldsymbol{\Theta}^{(j)})}.$
Within formula (13), the key part is $Pr(s_{i,t,h}=1,\boldsymbol{x}_{i}\mid\boldsymbol{z}_{i},\boldsymbol{\Theta}^{(j)})$. To calculate this probability, we adopt the idea of backward algorithm in dynamic programming (Rabiner, 1989). Then we give the definition (14) of backward probability $\boldsymbol{\beta}_{i,t,k}=(\beta_{i,t,k,h})_{h=1,\dots,M}^{\rm T}$, where
$\beta_{i,t,k,h}=Pr((x_{i,t+1},\dots,x_{i,T})^{\rm T},s_{i,t,h}=1\mid z_{i,k}=1,\boldsymbol{\Theta}^{(j)}).$ (14)
The value of $\boldsymbol{\beta}_{i,t,k}$ can be determined by iterative calculation (15).
@@
$\boldsymbol{\beta}_{i,t,k}$ $=$ $\boldsymbol{\Pi}_{k}^{\rm T}\cdot[(f(x_{i,t})\mid s_{i,t,h}=1,\boldsymbol{\Theta}^{(j)})_{h=1,\dots,M}\circ\boldsymbol{\beta}_{i,t+1,k}],\textnormal{when }t<T.$ (15)
Combine the forward probabilities and backward probabilities, we have the following equation to update (13),
$E(\boldsymbol{s}_{it}\mid\boldsymbol{z}_{i},\boldsymbol{x}_{i},\boldsymbol{\Theta}^{(j)})=Pr(s_{i,t},\boldsymbol{x}_{i}\mid\boldsymbol{z}_{i},\boldsymbol{\Theta}^{(j)})=\frac{\boldsymbol{\alpha}_{i,t,k}\circ\boldsymbol{\beta}_{i,t,k}}{\boldsymbol{\alpha}_{i,t,k}\cdot\boldsymbol{\beta}_{i,t,k}}.$
And then,
$E[I(s_{i,t,h}=1)]=\sum_{k=1}^{K}Pr(s_{i,t,h}=1\mid z_{i,k}=1,\boldsymbol{x}_{i},\boldsymbol{\Theta}^{(j)})Pr(z_{i,k}=1\mid\boldsymbol{x}_{i},\boldsymbol{\Theta}^{(j)}),$
where all values are calculated before. Following the notations from Rabiner (1989), we define $\gamma_{i,t,k,h}(\boldsymbol{\Theta}^{(j)})=E[I(s_{i,t,h}=1)\mid z_{i,k}=1,\boldsymbol{x}_{i},\boldsymbol{\Theta}^{(j)}]$ and $\eta_{i,t,h}(\boldsymbol{\Theta}^{(j)})=E[I(s_{i,t,h}=1)\mid\boldsymbol{x}_{i},\boldsymbol{\Theta}^{(j)}]$.
@@
$E[I(s_{i,t-1,p}=1,s_{i,t,q}=1)\mid\boldsymbol{z}_{i},\boldsymbol{x}_{i},\boldsymbol{\Theta}^{(j)}]$ (16)
$=$ $Pr[s_{i,t-1,p}=1,s_{i,t,q}=1\mid\boldsymbol{z}_{i},\boldsymbol{x}_{i},\boldsymbol{\Theta}^{(j)}]$
@@
We define $\xi_{i,t,k,p,q}(\boldsymbol{\Theta}^{(j)})=E[I(s_{i,t-1,p}=1,s_{i,t,q})=1\mid z_{i,k}=1,\boldsymbol{x}_{i},\boldsymbol{\Theta}^{(j)}]$.

### v2

@@ 5.2 Results from 3-Cluster MHMM-$\beta$ Model
for $h=1,\ldots,3$. As seen from both Figure 5 and the estimated parameters in Table 6, the three states clearly correspond to relatively low, medium, and high intensity levels. The low-intensity state (State 1) accommodates intensities in the range of about 0.00 to 0.50, with mean 0.290 and relatively the highest variance. The medium-intensity state (State 2) mostly concentrates within the 0.50 to 0.80 spectrum, with mean 0.597 and a relatively small variance. The high-intensity state (State 3) covers the range of about 0.75 to 1.00, with mean 0.787 and relatively the smallest variance. The zero and one inflation rates are also reflected in Figure 5. The low-intensity state accommodates the zero intensities, while the high-intensity state encompasses all the intensities elevated to one. We remark that because the standardized data contain virtually no “1” observations in State 1 and no “0” observations in States 2 and 3, the corresponding MLEs for $\epsilon_{h}^{(1)}$ (State 1) and $\epsilon_{h}^{(0)}$ (States 2-3) lie on the boundary, and the resulting asymptotic standard errors therefore appear extremely small.
@@
State 1 ($h=1$) State 2 ($h=2$) State 3 ($h=3$)
$\widehat{\epsilon}_{h}^{(0)}$ 0.025 $(4.917\times 10^{-3})$ 0.000 $(1.276\times 10^{-10})$ 0.000 $(6.301\times 10^{-15})$
$\widehat{\epsilon}_{h}^{(1)}$ 0.000 $(4.439\times 10^{-9})$ 0.001 $(2.714\times 10^{-3})$ 0.017 $(9.022\times 10^{-4})$
$\widehat{a}_{h}$ 2.195 $(8.580\times 10^{-3})$ 10.077 $(3.336\times 10^{-2})$ 11.658 $(1.803\times 10^{-3})$
$\widehat{b}_{h}$ 5.183 $(4.178\times 10^{-3})$ 6.805 $(9.739\times 10^{-4})$ 3.227 $(1.015\times 10^{-2})$
$\widehat{\mu}_{h}$ 0.290 0.597 0.787
$\widehat{\sigma}_{h}^{2}$ 0.0266 0.0137 0.0113
@@ 5.3 State Transition Patterns and Clusters of Quantum Dots
$\widehat{\boldsymbol{\pi}}_{1}=\begin{bmatrix}0.213&0.443&0.344\end{bmatrix},\widehat{\boldsymbol{\pi}}_{2}=\begin{bmatrix}0.000&0.143&0.857\end{bmatrix},\widehat{\boldsymbol{\pi}}_{3}=\begin{bmatrix}0.104&0.461&0.435\end{bmatrix}.$
Based on the estimated MHMM-$\beta$ model and according to the Bayes rule, there are 28, 54, and 46 QDs in Clusters 1, 2, and 3, respectively. To visualize, we randomly pick 3 dots from each cluster and presented their intensity series along with their estimated state series in Figures 6, 7, and 8, respectively.
@@
In Cluster 2, the QDs exhibit extremely high probabilities of staying in the medium or high-intensity states, and more importantly, the probabilities of transiting into the low-intensity state are essentially zero. This results in a stationary distribution showing that these QDs spend most of their time in the high-intensity state with probability 91.1%, a much smaller amount of time in the medium-intensity state with probability 8.9%, and ultimately no time at all in the low-intensity state. In Figure 7, the three randomly selected QDs from this cluster all stayed in the high-intensity state throughout the continuous excitation.
@@ 6 Conclusion & Discussion
This study introduces a novel statistical approach for analyzing the intensity fluctuation patterns of colloidal quantum dots (QDs), a critical problem in the fields of chemistry and materials science. The main challenge lies in the inherent complexity of the data: the intensity measurements exhibit intricate patterns influenced by unobserved states and are subject to substantial variation across individual QDs. Through a close collaboration with chemistry researchers, our proposed analytical pipeline and the mixture hidden Markov model with an inflated Beta distribution (MHMM-$\beta$) address these challenges by simultaneously modeling the intensity fluctuations and clustering the QDs based on their transition dynamics. This integrative approach captures the shared structure among QDs while allowing for individual differences, providing a more comprehensive understanding of their individual and collective behaviors under continuous excitation.
The findings from our analysis highlight the method’s effectiveness and its ability to generate interpretable results that align with principles of chemistry and physics. By identifying distinct clusters of QDs, each characterized by unique transition dynamics and stationary distributions, our approach provides chemists with valuable insights into the interplay between QD intensity states and their fluctuation styles.
There are several promising directions for future research. One critical and promising task is the joint analysis of photon lifetime and intensity. Recent chemical studies have emphasized the importance of understanding the relationship between photon lifetime, which reflects aspects of the QD microenvironment, and intensity fluctuations, which capture macroenvironmental behaviors. Developing methods that incorporate photon lifetime into the integrative analysis would enable us to reveal dependencies and connections between these two dimensions, offering a more holistic view of QD properties. Further methodological advancements could focus on developing more efficient algorithms for computation and implementing more accurate methods for model selection. From a broader perspective, the proposed approach could be adapted to address other scientific problems involving complex temporal data with state-dependent dynamics. For example, similar methods could be applied in neuroimaging to study brain activity patterns or in ecology to analyze animal movement behaviors.
@@ Acknowledgment
Appendix A Details on the EM Algorithm
@@ A.1 Derivation of the likelihood functions
$\begin{split}l(\boldsymbol{\Theta};\boldsymbol{X},\boldsymbol{S},\boldsymbol{Z})=&\sum_{i=1}^{N}\log\{\prod_{k=1}^{K}[\delta_{k}\boldsymbol{\pi}_{k,\boldsymbol{s}_{i,1}}f(x_{i,1}\mid\boldsymbol{s}_{i,1})\prod_{t=2}^{T}(\boldsymbol{\Pi}_{k})_{\boldsymbol{s}_{i,t-1},\boldsymbol{s}_{i,t}}f(x_{i,t}\mid\boldsymbol{s}_{i,t})]^{I(z_{i,k}=1)}\}\\ =&\sum_{i=1}^{N}\sum_{k=1}^{K}I(z_{i,k}=1)\log(\delta_{k})+\sum_{i=1}^{N}\sum_{k=1}^{K}I(z_{i,k}=1)\sum_{p=1}^{M}I(s_{i,1,p}=1)\log(\boldsymbol{\pi}_{k,p})\\ &+\sum_{i=1}^{N}\sum_{k=1}^{K}I(z_{i,k}=1)\log(\prod_{t=2}^{T}(\boldsymbol{\Pi}_{k})_{\boldsymbol{s}_{i,t-1},\boldsymbol{s}_{i,t}})\\ &+\sum_{i=1}^{N}\sum_{k=1}^{K}I(z_{i,k}=1)\log(\prod_{t=1}^{T}f(x_{i,t}\mid\boldsymbol{s}_{i,t}))\\ =&\sum_{i=1}^{N}\sum_{k=1}^{K}I(z_{i,k}=1)\log(\delta_{k})+\sum_{i=1}^{N}\sum_{k=1}^{K}I(z_{i,k}=1)\sum_{p=1}^{M}I(s_{i,1,p}=1)\log(\boldsymbol{\pi}_{k,p})\\ &+\sum_{i=1}^{N}\sum_{k=1}^{K}I(z_{i,k}=1)\{\sum_{t=2}^{T}\sum_{p=1}^{M}\sum_{q=1}^{M}I(s_{i,t-1,p}=1,s_{i,t,q}=1)\log((\boldsymbol{\Pi}_{k})_{p,q})\}\\ &+\sum_{i=1}^{N}\sum_{k=1}^{K}I(z_{i,k}=1)\{\sum_{t=1}^{T}\sum_{p=1}^{M}I(s_{i,t,p}=1)\log f(x_{i,t}\mid s_{i,t,p}=1)\}.\end{split}$
@@ A.2 E-Step
Let $\boldsymbol{\Theta}^{(j)}$ denote the set of parameter estimates from $j^{th}$ iteration. The expected complete-data log-likelihood function is then given by:
@@
In (7), we rewrite the integration w.r.t. the joint distribution $(\boldsymbol{S},\boldsymbol{Z})$ as a sequence of expectations, which could simplify the computation.
@@
$E[I(z_{i,k}=1)\mid\boldsymbol{x}_{i},\boldsymbol{\Theta}^{(j)}]$ $=$ $P(z_{i,k}=1\mid\boldsymbol{x}_{i},\boldsymbol{\Theta}^{(j)})$ (8)
$=$ $\frac{P(z_{i,k}=1,\boldsymbol{x}_{i}\mid\boldsymbol{\Theta}^{(j)})}{\sum_{k^{\prime}=1,\dots,K}P(z_{i,k^{\prime}}=1,\boldsymbol{x}_{i}\mid\boldsymbol{\Theta}^{(j)})}.$
In (8), the key part is $P(z_{i,k}=1,\boldsymbol{x}_{i}\mid\boldsymbol{\Theta}^{(j)})$. To calculate this probability, we adopt the idea of forward algorithm in dynamic programming (Rabiner, 1989). That is, the forward probabilities are defined as
$\alpha_{i,t,k,h}=P((x_{i,1},\dots,x_{i,t})^{\rm T},s_{i,t,h}=1\mid z_{i,k}=1,\boldsymbol{\Theta}^{(j)}).$
The value of $\boldsymbol{\alpha}_{i,t,k}$ can be determined through iterative computation as follows.
@@
$\boldsymbol{\alpha}_{i,t,k}$ $=$ $(\boldsymbol{\Pi}_{k}^{\rm T}\cdot\boldsymbol{\alpha}_{i,t-1,k})\circ(f(x_{i,t})\mid s_{i,t,h}=1,\boldsymbol{\Theta}^{(j)})_{h=1,\dots,M}^{\rm T},\textnormal{when }t>1,$
@@
Using the forward probabilities at the last time point $T$, we have that $P(z_{i,k}=1,\boldsymbol{x}_{i}\mid\boldsymbol{\Theta}^{(j)})=\lVert\boldsymbol{\alpha}_{i,T,k}\rVert_{1}=\sum_{h=1}^{M}\alpha_{i,T,k,h}$ and
$\tau_{i,k}(\boldsymbol{\Theta}^{(j)})\equiv E[I(z_{i,k}=1)\mid\boldsymbol{x}_{i},\boldsymbol{\Theta}^{(j)}]$ $=$ $\frac{\lVert\boldsymbol{\alpha}_{i,T,k}\rVert_{1}}{\sum_{k^{\prime}=1,\dots,K}\lVert\boldsymbol{\alpha}_{i,T,k^{\prime}}\rVert_{1}}$ (9)
@@
Here, following Rabiner (1989), we have denoted $E[I(z_{i,k}=1)\mid\boldsymbol{x}_{i},\boldsymbol{\Theta}^{(j)}]$ as $\tau_{i,k}(\boldsymbol{\Theta}^{(j)})$.
After evaluating the inner expectation, the outer expectation is computed as follows:
(ii) Compute $I(s_{i,t,h}=1)$ :
Only $E[I(s_{i,t,h}=1)]$ is explicitly solved and $E[I(s_{i,1,h}=1)]$ is treated as a special case of the former.
$E[I(s_{i,t,h}=1)\mid\boldsymbol{z}_{i},\boldsymbol{x}_{i},\boldsymbol{\Theta}^{(j)}]$ $=$ $P(s_{i,t,h}=1\mid\boldsymbol{z}_{i},\boldsymbol{x}_{i},\boldsymbol{\Theta}^{(j)})$ (10)
$=$ $\frac{P(s_{i,t,h}=1,\boldsymbol{x}_{i}\mid\boldsymbol{z}_{i},\boldsymbol{\Theta}^{(j)})}{\sum_{h^{\prime}=1,\dots,M}P(s_{i,t,h^{\prime}}=1,\boldsymbol{x}_{i}\mid\boldsymbol{z}_{i},\boldsymbol{\Theta}^{(j)})}.$
In (10), the key part is $P(s_{i,t,h}=1,\boldsymbol{x}_{i}\mid\boldsymbol{z}_{i},\boldsymbol{\Theta}^{(j)})$. To calculate this probability, we adopt the idea of backward algorithm in dynamic programming (Rabiner, 1989). We write $\boldsymbol{\beta}_{i,t,k}=(\beta_{i,t,k,h})_{h=1,\dots,M}^{\rm T}$, where
$\beta_{i,t,k,h}=P((x_{i,t+1},\dots,x_{i,T})^{\rm T},s_{i,t,h}=1\mid z_{i,k}=1,\boldsymbol{\Theta}^{(j)}).$
The value of $\boldsymbol{\beta}_{i,t,k}$ can be determined by iterative calculations as follows.
@@
$\boldsymbol{\beta}_{i,t,k}$ $=$ $\boldsymbol{\Pi}_{k}^{\rm T}\cdot[(f(x_{i,t})\mid s_{i,t,h}=1,\boldsymbol{\Theta}^{(j)})_{h=1,\dots,M}\circ\boldsymbol{\beta}_{i,t+1,k}],\textnormal{when }t<T.$
Combine the forward probabilities and backward probabilities, (10) can be updated as:
$\gamma_{i,t,k,h}(\boldsymbol{\Theta}^{(j)})$$\equiv$ $E[I(s_{i,t,h}=1)\mid z_{i,k}=1,\boldsymbol{x}_{i},\boldsymbol{\Theta}^{(j)}]=P(s_{i,t,h}=1\mid z_{i,k}=1,\boldsymbol{x}_{i},\boldsymbol{\Theta}^{(j)})$
$=$ $\frac{\boldsymbol{\alpha}_{i,t,k}\circ\boldsymbol{\beta}_{i,t,k}}{\boldsymbol{\alpha}_{i,t,k}^{\rm T}\cdot\boldsymbol{\beta}_{i,t,k}},$
and
$\eta_{i,t,h}(\boldsymbol{\Theta}^{(j)})$$\equiv$ $E[I(s_{i,t,h}=1)\mid\boldsymbol{x}_{i},\boldsymbol{\Theta}^{(j)}]=\sum_{k=1}^{K}P(s_{i,t,h}=1\mid z_{i,k}=1,\boldsymbol{x}_{i},\boldsymbol{\Theta}^{(j)})P(z_{i,k}=1\mid\boldsymbol{x}_{i},\boldsymbol{\Theta}^{(j)})$
$=$ $\sum_{k=1}^{K}\gamma_{i,t,k,h}(\boldsymbol{\Theta}^{(j)})\tau_{i,k}(\boldsymbol{\Theta}^{(j)}).$
We use $\gamma_{i,t,k,h}(\boldsymbol{\Theta}^{(j)})$ to denote $E[I(s_{i,t,h}=1)\mid z_{i,k}=1,\boldsymbol{x}_{i},\boldsymbol{\Theta}^{(j)}]$ and use $\eta_{i,t,h}(\boldsymbol{\Theta}^{(j)})$ to denote $E[I(s_{i,t,h}=1)\mid\boldsymbol{x}_{i},\boldsymbol{\Theta}^{(j)}]$.
@@
$\xi_{i,t,k,p,q}(\boldsymbol{\Theta}^{(j)})$ $=$ $E[I(s_{i,t-1,p}=1,s_{i,t,q}=1)\mid\boldsymbol{z}_{i},\boldsymbol{x}_{i},\boldsymbol{\Theta}^{(j)}]$
$=$ $P[s_{i,t-1,p}=1,s_{i,t,q}=1\mid\boldsymbol{z}_{i},\boldsymbol{x}_{i},\boldsymbol{\Theta}^{(j)}]$
@@
We denote $E[I(s_{i,t-1,p}=1,s_{i,t,q}=1)\mid z_{i,k}=1,\boldsymbol{x}_{i},\boldsymbol{\Theta}^{(j)}]$ as $\xi_{i,t,k,p,q}(\boldsymbol{\Theta}^{(j)})$.